kristen00001:

Write the equation of a line that has a slope of -1 and passes through the point (-4,9). Describe/show your work.

Vocaloid:

Point slope form y - y1 = m(x - x1) m is the slope, which is given as m = -1 (x1, y1) is your point, which is given as (-4,9), so your x1 = -4 and y1 = 9 Plug in the appropriate quantities. If your equation must be in y = mx + b form, simply distribute the m and solve for y

kristen00001:

thanks